OpenStudy (anonymous):

f(x)= 2x+5x f(3) ??? solve this.. :)

OpenStudy (anonymous):

if f is = to 3

mathslover (mathslover):

@mjay , put x = 3 to solve for f(3) If f(x) = 2 + x then , f(2) = 2 + 2 = 4 similarly here f(x) = 2x + 5x then f(3) = 2(3) + 5(3)

OpenStudy (anonymous):

21

OpenStudy (anonymous):

no it's 10

OpenStudy (anonymous):

its 21.

OpenStudy (anonymous):

why 21..???

OpenStudy (anonymous):

Because you plug in 3 for every x value there is in the function. f(x)= 2x+5x f(3)= 2(3)+5(3) =6+15 =21